A sample of size 3 is taken from a population of the number of students in each class of Grade 9 students listed as 25, 30, 35, and 40.





Construct the sampling distribution of the sample means






1
Expert's answer
2022-03-19T02:41:07-0400

Solution

Population size "N=4"

Sample size "n=3"


Possible samples

"=C_4^3= \binom{4}{3}=4" samples


"\def\arraystretch{1.5}\begin{array}{c:c:c} no.& sample &mean\\\hline1.&25,30,35&30.00\\\hdashline2.&25,30,40&31.67\\\hdashline3.&25,35,40&33.33\\\hdashline4.&30,35,40&35.00\\\hline\end{array}"


Probability distribution

"\def\arraystretch{1.5}\begin{array}{c:c:c} mean&freq&prob\\\hline30.00&1&1/4\\\hdashline31.67&1&1/4\\\hdashline33.33&1&1/4\\\hdashline35.00&1&1/4\\\hline\end{array}"


Histogram